Two triangles can be formed with the given information. Use law of sines to solve the triangle
B=46degrees , a=12 , b=11
1
Expert's answer
2020-03-10T13:26:06-0400

a/sinA=b/sinB=c/sinC

sinA=(a*sinB) /b=(12*sin46°) /11=0.78

Variant 1

A=arcsin(0.78)=51°,

C=180°-46°-51°=83°,

c=(b*sinC)/sinB=(11*sin83°)/sin46°=15.2


Variant 2

A=180°-51°=129°,

C=180°-46°-129°=5°,

c=(b*sinC)/sinB=(11*sin5°)/sin46°=1.33